# Exploiting Kubernetes’ Image Pull Implementation to Deny Node Availability

### v1

@@ Abstract
Kubernetes (K8s) has grown in popularity over the past few years to become the de-facto standard for container orchestration in cloud-native environments. While research is not new to topics such as containerization and access control security, the Application Programming Interface (API) interactions between K8s and its runtime interfaces have not been studied thoroughly. In particular, the CRI-API is responsible for abstracting the container runtime, managing the creation and lifecycle of containers along with the downloads of the respective images. However, this decoupling of concerns and the abstraction of the container runtime renders K8s unaware of the status of the downloading process of the container images, obstructing the monitoring of the resources allocated to such process. In this paper, we discuss how this lack of status information can be exploited as a Denial of Service attack in a K8s cluster. We show that such attacks can generate up to 95% average CPU usage, prevent downloading new container images, and increase I/O and network usage for a potentially unlimited amount of time. Finally, we propose two possible mitigation strategies: one, implemented as a stopgap solution, and another, requiring more radical architectural changes in the relationship between K8s and the CRI-API.
@@ I Introduction
In particular, when deploying a Pod— a group of one or more containers — an image pull is usually triggered for the corresponding containers if unavailable locally. If a container is deleted before the image is fully downloaded, the API will not communicate this action to the container runtime. However, the runtime will still download the image, wasting resources in the process. Similarly, when a container is forcibly destroyed, its resources are not instantly released. While K8s will delete all information from the API as soon it receives the request, the complete elimination of the containers composing the Pod will only occur when all modules finish their deletion processes. Although this is intended behavior, such a process could be exploited to mount a Denial of Service (DoS) attack on the cluster. Indeed, these image download requests would prevent other users from scheduling new jobs without a cached image and force the tenant to consume unnecessary resources.
@@ II-A Kubernetes
Kubernetes is an open-source orchestration system for automating deployment, scaling, and management of containerized applications, initially developed by Google and maintained by the Cloud Native Computing Foundation (CNCF) [8]. K8s is built from a set of composable modules through a set of standard APIs that can be extended by the users alongside the core components [1]. Over time, this allowed the development of several new infrastructure paradigms, like Function as a Service (FaaS) frameworks[9] and multi-cluster management [6].
@@ III-A Threat Model and Assumptions
As described in the Role-Based Access Control (RBAC) Best Practices official documentation [^2] https://kubernetes.io/docs/concepts/security/rbac-good-practices/, it is expected that a single cluster will have several distinct SAs to limit the access to specific namespaces, to share the cluster between production and development, or to enable a continuous integration/delivery solution. So, it is reasonable to assume that one of these service accounts could be leaked from a development machine, a non-protect staging deployment used as a vector to access an inside container with a mount SA token, or a insider malicious user with access to the cluster.
@@
On the restriction side, we understand that a compromised container or a SA token could be used for more advanced attacks, such as lateral movement [7], Out-of-Memory (OOM) attack [12], or deploying cryptomining in the node [13]. However, all these attacks can be straightforwardly identified by inspecting the resource usage or checking the elements running in the cluster. Notwithstanding, this attack wants to create a stealthy channel to cripple the cluster’s resources that cannot be easily identifiable without an extensive evaluation of the node resources and processes, differing it from the other ones.
On the other hand, the attack can be performed with the same minimal permissions on a wide variety of settings, ranging from busy clusters with several concurrent deployments to smaller, soft multitenancy clusters [^3] Soft-multitenancy refer to resource sharing and strong trust between the tenants. https://kubernetes.io/docs/concepts/security/multi-tenancy/#isolation, and multi-cluster topologies [14]. Even with monitoring or auditing logs active, we argue that the attack will be hard to discover without a specific and deeper analysis of the node itself, which is not a trivial task. Indeed, even in cloud deployments — such as Google Kubernetes Engine (GKE), Azure Kubernetes Service (AKS), and Amazon Elastic Kubernetes Service (EKS) — monitoring tools tend to provide an aggregated view of the cluster’s resource usage, and administrators usually will not access directly the cluster nodes.
We further assume that the node and guest Pods have no known vulnerabilities and that all security mechanisms work properly.
@@ III-B Asynchronous API Communication Status
As described in subsection II-C, K8s uses an asynchronous communication to enforce the actions in the several interfaces implemented in the cluster. For example, each Pod creation is executed independently by K8s, but each node’s containerd will have the last say in defining the order of execution for each request.
@@
Since after sending the request K8s maintains no control over the request execution, it relies on the interface implementation to update the intermediate and final status on the object. However, this does not always happen. The CRI has no API related to the status of image download, and so, the only response containerd will provide will be either the download confirmation or some error on the downloading process, like a missing secret or an unavailable tag.
The Kubernetes Enhancement Proposals (KEP) 3542444GitHub issue available at https://github.com/kubernetes/enhancements/issues/3542 proposes an update to the Kubelet, enabling download status updates to the API, solving the problem previously described. Even with this KEP, the problem is not entirely solved: indeed, it remains impossible to cancel pending downloads.
III-C Motivating Example
The attacker aims to select one or more worker nodes within the Kubernetes cluster and starve their resources, eventually causing a resource DoS condition. Once the attack has been successfully mounted, nodes can no longer schedule new Pods whose images are not already present in the local cache. Besides that, the CPU, disk, and network usage will also dramatically increase.
To attain this, attackers can exploit the aforementioned lack of status updating in CRI for image downloads. The attack strategy is simple: the attacker sends a Pod creation request, waits for the API to register that the image needs to be downloaded, and then deletes the Pod. This will not interrupt the download that has already been started. Repeating this multiple times will cause a substantial, sustained increase in resource consumption.
As said before, this attack is hard to identify and mitigate, since the requests are legitimate, and could be easily confused with a normal user making some API requests. On the other hand, the node will suffer from an high and possibly irregular workload, as each image is downloaded and unpacked. Furthermore, the inability to download images may be attributed to an external issue with the container registry itself, and not to a local DoS attack.
@@ IV Attack Description
The following section proposes two attacks that exploit the aforementioned vulnerability in containerd and the CRI-API, first rendering the Kubernetes node unable to pull any Pod image for an extended period and, secondly, impacting a node’s cache, increasing the amount of time needed to deploy an instance of an application.
@@ IV-A Attack Vector
While this choice has its roots in the separation of concerns between Kubernetes and containerd, it has a subtle side effect. Without having to notify the API about the status of the image pull, containerd is completely unaware of the addition or removal of Pods through the Kubernetes API. With that, a malicious actor can try to use a significant amount of resources by repeatedly instantiating new Pods in the node. However, K8s addresses this with the implementation of Quotas, which limit the number of objects that can be created in a namespace. The malicious actor can overcome this limitation by deleting objects immediately after the instantiation. Nevertheless, Quotas will still apply until the object is fully removed from the cluster.
@@
As discussed in subsection II-C, one of the ways to immediately release resources is to use the ---force flag. This will automatically remove the API bindings for the resources. On the other hand, it will not interrupt any image download that has already been started by containerd. [^5] At the time of writing, a bug exists in K8s that additionally prevents the deletion of associated containers when a Pod is force deleted. See https://github.com/kubernetes/kubernetes/issues/119276. Thus, after deploying a Pod an attacker can force delete it shortly after, essentially bypassing Quotas and avoiding having to wait for the image downloads to complete. This process can be repeated a potentially infinite number of times, in quick succession. By doing so, an attacker can clutter the containerd’s download queue with useless image pull requests as long as they have image tags available to pull. Since containerd is agnostic to the Pods’ existence and Kubernetes is unaware about what images are being downloaded, the status quo can remain indefinitely. As such, the target container runtime will continuously be forced to keep track of and download useless container images on behalf of the attacker.
IV-A1 Attack Requirements
@@
The attack requires minimal prerequisites to be executed. First, it needs access to a cluster with standard user privileges and the capacity to create and delete at least one Pod. Any limitation imposed by the RBAC configuration, Quotas, and Pod Security Standards is irrelevant to its execution. Second, the attacker needs a set of Docker images, with different tags of a single image, or entirely different ones altogether. They also need to be in a registry accessible by the node. Furthermore, the least layers the images have in common, the better, since re-occurring layers will not be re-downloaded.
@@
The content, structure, functionality, and size of the image are irrelevant for the attack to be successful. Their only limiting aspect is that the total size of the image cannot exceed eventual limits imposed by the cluster administrator because the download will be blocked before starting.
@@
IV-B Denial of Resources
This attack will generate a denial of resource in the node, and its impact is twofold. First, the endless cycle of image pull requests and deletions prevents the deployment of new containers without a cached image. While this impacts less long-running tasks such as web services, it prevents short-lived workloads, such as Continuous Integration/Continuous Delivery (CI/CD) pipelines, from deploying and completing in a few seconds or minutes, causing a loss of availability. This problem is exacerbated by the growing trend of running short-lived microservices: indeed, 72% of containers live fewer than five minutes [15]. An unresponsive cluster can be extremely problematic, even if the attack is short lived: clusters that rely on short and fast batch jobs (e.g. CI/CD) will be inevitably delayed.
Second, the attack drains the available resources on the worker node by forcing it to constantly utilize disk, bandwidth, and CPU to retrieve container images. As a result, other workloads hosted on the node may suffer a performance setback, possibly violating Service Level Agreements. Since the Kubelet usually runs as a process in the node and it is not directly controlled by the quotas and limitation imposed by the Kubernetes, it will use 100% of the resources available to download and decompress the images, and by default, there is no way to limit this resource access.
IV-C Cache Manipulation
@@
IV-D Responsible Disclosure
The attack presented in this work has been responsibly disclosed to the Kubernetes Security Team. They informed us that DoS via Pod creation is considered a resource usage bug. Nevertheless, we are actively collaborating with them to address this issue, and we have opened a public issue  [^7] The link for the issue to solve this in future versions.
@@ V-A Experimental Setup
We set up a local testbed on our premises to evaluate the attack. On a machine equipped with an Intel Xeon Silver 4112 CPU (2.60GHz), we created a virtualized environment with a master and two worker nodes without any special software configuration. Each machine has 2 vCPUs, 4 GB of memory, 120 GB of storage capacity (HDD), a 1 Gbit/s network interface, and Ubuntu 22.04 LTS as its Operating System (OS).
While to the best of our knowledge no information about the average size of K8s nodes can be found in literature, we tried to match the specifications of our nodes to the default one that users find when creating a K8s cluster on GKE [^8] GKE uses a e2-medium machine type with 2 shared vCPUs, 4 GB of memory and 100 GB of storage capacity per node., Linode999www.linode.com/pricing/ or DigitalOcean101010www.digitalocean.com/pricing/kubernetes. On the other hand, AKS does not provide default values but suggests slightly higher ones in their best practices [^11] https://github.com/Azure/k8s-best-practices/blob/master/Cost_Optimization.md#node---vm-sizes.
Finally, we used a separate physical machine to create an additional Virtual Machine (VM) acting as a private Docker registry. The registry and the cluster also communicate through a 1 Gbit/s network interface.
To further confirm our findings, we additionally deploy a cluster on GKE, using n2d-standard-2 machines with 2 AMD Epyc vCPUs, 8 GB of RAM, a 100 GB balanced persistent disk and a 10 Gbit/s network interface. All the experiments outlined in this section have been successfully replicated in both the public and private cloud scenarios, with comparable results. Due to space constraints and unless specifically mentioned, the following results and graphs refer to the private cloud experiments.
V-A1 Images
@@
By researching the Docker image compression mechanism, which relies on gzip, we discover that different types of randomly generated data influence the decompression time of archives. For example, random English text is efficiently and quickly compressed by gzip, while random data formatted with the od command yields varying times and compressing rates. The details of this research are available on GitHub [16].
Exploiting these results, we create two sets of images to assess the attack’s effectiveness on two distinct image sizes: one with large layers focused on bandwidth usage and the second with small layers focused on the I/O overhead.
Each image in the set contains a base Ubuntu layer (of around 80 MB uncompressed), which is shared between all the images, and a variable number of layers, created with the od command and the c option. Moreover, by forcing the --no-cache flag, we make sure that every layer is always re-built and, therefore, unique in terms of content.
@@
Lower-bound uncompressed size 2.12GB 0.10GB
Upper-bound uncompressed size 14.40GB 0.90GB
Total uncompressed size* 57.12GB 16.80GB
Total compressed size* 28.88GB 8.47GB

### v2

@@ Abstract
Kubernetes (K8s) has grown in popularity over the past few years to become the de-facto standard for container orchestration in cloud-native environments. While research is not new to topics such as containerization and access control security, the Application Programming Interface (API) interactions between K8s and its runtime interfaces have not been studied thoroughly. In particular, the CRI-API is responsible for abstracting the container runtime, managing the creation and lifecycle of containers along with the downloads of the respective images. However, this decoupling of concerns and the abstraction of the container runtime renders K8s unaware of the status of the downloading process of the container images, obstructing the monitoring of the resources allocated to such process. In this paper, we discuss how this lack of status information can be exploited as a Denial of Service attack in a K8s cluster. We show how such attacks can impact worker nodes, generating up to 95% average CPU usage, prevent downloads of new container images, and increase I/O and network usage for a potentially unlimited amount of time. We argue that solving this problem would require a radical architectural change in the relationship between K8s and the CRI-API, which would be unfeasible in the short term. Thus, as a stopgap solution, we propose MAGI: an eBPF-based, proof-of-concept mitigation that detects and terminates potential attacks.
@@ I Introduction
In particular, when deploying a Pod— a group of one or more containers — an image pull is usually triggered for the corresponding containers if unavailable locally. If a container is deleted before the image is fully downloaded, the API will not communicate this action to the container runtime. However, the runtime will still download the image, wasting resources in the process. Similarly, when a container is forcibly destroyed, its resources are not instantly released. While K8s will delete all information from the API as soon it receives the request, the complete elimination of the containers composing the Pod will only occur when all modules finish their deletion processes. Although this is intended behavior, such a process could be exploited to mount a Denial of Service (DoS) attack on the cluster’s worker nodes. Indeed, these image download requests would prevent other users from scheduling new jobs without a cached image and force the tenant to consume unnecessary resources.
@@ II-A Kubernetes
Kubernetes is an open-source orchestration system for automating deployment, scaling, and management of containerized applications, initially developed by Google and maintained by the Cloud Native Computing Foundation (CNCF[8]. K8s is built from a set of composable modules through a set of standard APIs that can be extended by the users alongside the core components [1]. Over time, this allowed the development of several new infrastructure paradigms, like Function as a Service (FaaS) frameworks [9] and multi-cluster management [6].
@@ III-A Threat Model and Assumptions
As described in the Role-Based Access Control (RBAC) Best Practices official documentation [^2] https://kubernetes.io/docs/concepts/security/rbac-good-practices/, it is expected that a single cluster will have several distinct SAs to limit the access to specific namespaces, to share the cluster between production and development, or to enable a continuous integration/delivery solution. So, it is reasonable to assume that one of these service accounts could be leaked from a development machine, a non-protected staging deployment used as a vector to access an inside container with a mount SA token, or an insider malicious user with access to the cluster.
@@
On the restriction side, we understand that a compromised container or an SA token could be used for more advanced attacks, such as lateral movement [7], Out-of-Memory (OOM) attack [12], or deploying cryptomining in the node [13]. However, all these attacks can be straightforwardly identified by inspecting the resource usage or checking the elements running in the cluster. Notwithstanding, this attack wants to create a stealthy channel to cripple the cluster’s resources that cannot be easily identifiable without an extensive evaluation of the node resources and processes, differentiating it from the other ones.
On the other hand, the attack can be performed with the same minimal permissions on a wide variety of settings, ranging from busy clusters with several concurrent deployments to smaller, soft multitenancy clusters [^3] Soft-multitenancy refers to resource sharing and strong trust between the tenants. https://kubernetes.io/docs/concepts/security/multi-tenancy/#isolation, and multi-cluster topologies [14]. Attackers can choose to target a specific node, for example, to target a GPU-enabled node. Using a nodeSelector or podAffinity, attackers can ensure that the Pod are scheduled on the desired node(s), further increasing the attack’s impact. Even with monitoring or auditing logs active, we argue that the attack will be hard to discover without a specific and deeper analysis of the node itself, which is not a trivial task. Indeed, even in cloud deployments — such as Google Kubernetes Engine (GKE), Azure Kubernetes Service (AKS), and Amazon Elastic Kubernetes Service (EKS) — monitoring tools tend to provide an aggregated view of the cluster’s resource usage, and administrators usually will not access directly the cluster nodes.
The registry employed by the attacker can be any public or private registry, as long as it is reachable from the cluster. Some public registries, like DockerHub, provide strict rate limits, which are computed per IP. Since the IP pulling the image is the node’s IP, being banned translates into a denial of service for the node itself, which is exactly what the attacker wants. On the other hand, a private registry can also be used, allowing attackers to have a near-infinite arsenal of images to download.
Finally, we assume that the node and guest Pods have no known vulnerabilities and that all security mechanisms work properly.
@@ III-B Asynchronous API Communication Status
As described in subsection II-C, K8s uses an asynchronous communication to enforce the actions in several interfaces implemented in the cluster. For example, each Pod creation is executed independently by K8s, but each node’s containerd will have the last say in defining the order of execution for each request [^4] The CRI specification does not mandate a specific logic for handling the order in which image pull requests are served, leading to possible differences between container runtimes..
@@
Since after sending the request, K8s maintains no control over the request execution, it relies on the interface implementation to update the intermediate and final status on the object. However, this does not always happen. The CRI has no API related to the status of image download, and so, the only response containerd will provide will be either the download confirmation or some error on the downloading process, like a missing secret or an unavailable tag.
The Kubernetes Enhancement Proposals (KEP) 3542555GitHub issue available at https://github.com/kubernetes/enhancements/issues/3542 proposes an update to the Kubelet, enabling download status updates to the API, solving the problem previously described. Even with this KEP, the problem is not entirely solved: indeed, it remains impossible to cancel pending downloads. Even using alternative container runtimes such as CRI-O or more secure ones like gVisor or Kata Containers does not change the situation. All these runtimes still rely on the CRI interface, which suffers from the issue described above.
@@ IV Attack Description
The following section proposes two attacks that exploit the aforementioned vulnerability in containerd and the CRI-API, first rendering a Kubernetes node unable to pull any Pod image for an extended period and, secondly, impacting a node’s cache, increasing the amount of time needed to deploy an instance of an application.
@@ IV-A Attack Vector
While this choice has its roots in the separation of concerns between Kubernetes and containerd, it has a subtle side effect. Without having to notify the API about the status of the image pull, containerd is completely unaware of the addition or removal of Pods through the Kubernetes API. With that, a malicious actor can try to use a significant amount of resources by repeatedly instantiating new Pods in the node.
To further strengthen the attack, attackers can use the ---force flag when deleting a Pod. As discussed in subsection II-C, the ---force flag automatically removes all API bindings for the resources, which means that the Pod will be deleted immediately from the cluster. On the other hand, it will not interrupt any image download that has already been started by containerd. [^6] At the time of writing, a bug exists in K8s that additionally prevents the deletion of associated containers when a Pod is force deleted. See https://github.com/kubernetes/kubernetes/issues/119276. The bug is not planned to be fixed..
@@
Thus, after deploying a Pod an attacker can force delete it shortly after, and avoiding a lengthy wait for the image downloads to complete. This process can be repeated a potentially infinite number of times, in quick succession. By doing so, an attacker can clutter the containerd’s download queue with useless image pull requests as long as they have image tags available to pull. Since containerd is agnostic to the Pods’ existence and Kubernetes is unaware about what images are being downloaded, the status quo can remain indefinitely. As such, the target container runtime will continuously be forced to keep track of and download useless container images on behalf of the attacker.
IV-B Attack Requirements
The attack requires minimal prerequisites to be executed. First, it needs access to a cluster with standard user privileges and the capacity to create and delete at least one Pod. Any limitations imposed by the RBAC configuration or Pod Security Standard (PSS) are irrelevant to its execution. Indeed, as the vulnerability sits in the CRI interface, any hardening or security measures applied to abstraction layers above the CRI will not prevent the attack.
Quotas are also easily bypassed: the minimum requirement is to be able to create and delete a single Pod in a single namespace. K8s’ Quotas are designed to limit the number of objects that can be created in a namespace by a single user, preventing resource exhaustion attacks and ensuring fair resource distribution among users. However, since the attack relies on the constant deletion and recreation of Pods, a Quota of just one Pod is sufficient to execute the attack.
@@ IV-B Attack Requirements
Second, the attacker requires a set of Docker images, which can be either different tags of a single image or entirely different ones altogether. They also need to be in a registry accessible by the node. Furthermore, the least layers the images have in common, the better, since re-occurring layers will not be re-downloaded.
@@
The content, structure, functionality, and size of the image are irrelevant for the attack to be successful. Their only limiting aspect is that the total size of the image cannot exceed eventual limits imposed by the cluster administrator because the download will be blocked before starting. Thus, attackers can use any image they have access to, including public images from Docker Hub or private images from their own registries, as long as they are accessible by the node. Such images either need not to be present on the target node, or their download must be forced using the imagePullPolicy: always parameter.
@@
IV-C Denial of Resources
This attack will generate a denial of resource in the node, and its impact is twofold. First, the endless cycle of image pull requests and deletions prevents the deployment of new containers without a cached image. This problem is exacerbated by the growing trend of running short-lived microservices: indeed, 72% of containers live fewer than five minutes [15]. An unresponsive cluster can be extremely problematic, even if the attack is short-lived: clusters that rely on short and fast batch jobs (e.g. Continuous Integration/Continuous Delivery (CI/CD)) will inevitably be delayed. Indeed, this DoS does not render a node actually unavailable for K8s: the scheduler will still see the node as healthy and will continue to schedule Pods on it, nor any Pod on it will be evicted.
Second, the attack drains the available resources on the worker node by forcing it to constantly utilize disk, bandwidth, and CPU to retrieve container images. As a result, other workloads hosted on the node may suffer a performance setback, possibly violating Service Level Agreements. Since the Kubelet usually runs as a process in the node and it is not directly controlled by the quotas and limitations imposed by the Kubernetes, it will use 100% of the resources available to download and decompress the images, and by default, there is no way to limit this resource access. Furthermore, pinpointing the resource starvation to the attack is not straightforward, as the containerd process will appear as a busy process in the node, and the API will not provide any information about the ongoing image downloads.
Ultimately, the effectiveness of the denial of service will depend on the amount of resources available in the node and the amount of nodes in the cluster. Generally, K8s worker nodes have a limited size, and the K8s philosophy promotes smaller but more numerous nodes, enabling horizontal autoscaling and a better failure tolerance. Additionally, K8s is increasingly deployed in Edge environments, where nodes are often resource-constrained and have no redundancy. In these scenarios, the attack can be particularly effective, as it can quickly exhaust the available resources. On the other hand, larger nodes with more resources can withstand a higher number of image pull requests, but the attack can still cause a noticeable performance degradation.
Nonetheless, the amount of nodes in the cluster also has a direct impact on the attack. When attackers target a single or a few nodes, it can be easily attributed to a local malfunction, and its resource usage will be hard to pinpoint in an aggregated dashboard. On the other hand, executing the attack on many or all nodes in big clusters will likely cause a noticeable increase in resource usage across the cluster, making it easier to identify.
Finally, the possibility of a DoS is increased in multi-tenant scenarios such as Liqo [16], where resources are shared across clusters. As an example, a cluster may lend a subset of its resources to another cluster, allowing the latter to run workloads on it. To the renting cluster, this shows up as a local, virtual namespace, while Pods are actually running remotely. In this case, an attacker can target the shared namespace, causing a denial of service on the remote node(s).
IV-D Cache Manipulation
@@
IV-E Responsible Disclosure
The attack presented in this work has been responsibly disclosed to the Kubernetes community. We opened a public issue to discuss the attack and its implications [^8] GitHub issue available at https://github.com/kubernetes/kubernetes/issues/122905. The issue was acknowledged by the Kubernetes maintainers, but as of the time of writing, a solution has not yet been implemented in the codebase. Still, the maintainers have confirmed that the attack is not a security vulnerability, but rather a resource usage bug, spurring a discussion and several pull requests on how to address it in future versions of Kubernetes.
@@ V-A Experimental Setup
To evaluate the attack, we conducted a series of experiments on a K8s cluster. The goal was to assess the attack’s impact on resource usage, scheduling delays, and service interference. We also wanted to verify the attack’s effectiveness in manipulating the local image cache, and finally, to confirm the attack’s applicability in public cloud environments.
V-A1 Local testbed
We set up a local testbed on our premises to evaluate the attack. On a machine equipped with an Intel Xeon Silver 4112 CPU (2.60GHz) and 64 GB of RAM, we created three virtualized nodes with 2 vCPUs, 4 GB of memory, 120 GB of storage capacity (HDD), a 1 Gbit/s network interface, and Ubuntu 22.04 LTS as the operating system, to run as nodes in a K8s cluster: one for the control plane and two as worker nodes, using K8s v1.27 and containerd v1.6.8 as the container runtime. We also used a separate physical machine hosting a private Docker registry.
V-A2 Public cloud testbed
To further confirm our findings in bigger and more realistic scenarios, we also conducted some experiments in a public cloud environment. We used Google Cloud Platform’s GKE service, allowing us to assess the attack’s impact on a larger scale and in a more production-like environment. We used e2-standard-2 machines with 2 dedicated AMD Epyc vCPUs, 8 GB of RAM, a 60 GB balanced persistent disk, and a 10 Gbit/s network interface. We deployed a total of ten worker nodes, while the control plane was provided by GKE. The nodes were configured to use Ubuntu 24.04 LTS as the operating system and containerd as the container runtime. No particular settings were enabled during the creation of the cluster.
V-A3 Image sets
@@
Each image in the set contains a base Ubuntu layer (of around 80 MB uncompressed), which is shared between all the images, and a variable number of layers, created with the od command and the c option. We performed some experiments on gzip’s compression options, and the results are available in a separate report [17].
@@
Lower-bound
uncompressed size
2.12GB 0.10GB
Upper-bound
uncompressed size
14.40GB 0.90GB
Total
uncompressed size*
57.12GB 16.80GB
Total
compressed size*
28.88GB 8.47GB